Investigate to see if (2x+5) is a factor of
2x^2-3x^2-12x+20=0
Please solve the equation

Answer by stanbon(75887)

Investigate to see if (2x+5) is a factor of
2x^3-3x^2-12x+20=0
Please solve the equation
-----
If 2x+5 is a factor, x= -5/2 is a zero.
Check using synthetic division:
-5/2)....2....-3....-12....20
..........2.....-8....8....|..0

Since the remainder is zero, 2x+5 is a factor.
===================
Cheers,
Stan H.